OpenStudy (anonymous):

graph the line passing through the given point and having the given slope. Give the slope-intercept form of the equation of the line if possible. (1,-3) m=-2/5 Slope intercept form of the equation of the line. Can anyone help?

OpenStudy (anonymous):

use the almighty point slope formula \[y-y_1=m(x-x_1)\] with \[x_1=1,y_1=-3, m =-\frac{2}{5}\]

OpenStudy (anonymous):

if you write it out exactly you get \[y-(-3)=-\frac{2}{5}(x-1)\]

OpenStudy (anonymous):

distributive law \[y+3=-\frac{2}{5}x+\frac{2}{5}\]

OpenStudy (anonymous):

and finally subtract 3 from both sides to get \[y=-\frac{2}{5}x-\frac{13}{5}\]

OpenStudy (anonymous):

Ok, how did you get 13/5. I am having the worst time figuring that part out!

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

yeah it is fractions

OpenStudy (anonymous):

first of all it is not \[\frac{13}{5}\] it is \[-\frac{13}{5}\]

OpenStudy (anonymous):

is it clear where the \[\frac{2}{5}\] came from? because the next step was just \[\frac{2}{5}-3=\frac{2}{5}-\frac{15}{5}=\frac{2-15}{5}=\frac{-13}{5}\]

OpenStudy (anonymous):

I'm not trying to aggravate you, I just have a hard time here. Why did it end up being -15/5?

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Can you help me with this last question?